OpenStudy (anonymous):

Which are measurements of the sides of a right triangle? 202,198, 40 28, 26, 12 17, 14, 6, all of the above?

jimthompson5910 (jim_thompson5910):

Use the formula a^2 + b^2 = c^2 where 'c' is the hypotenuse and the longest side

OpenStudy (anonymous):

I dont know how

jimthompson5910 (jim_thompson5910):

For instance, in the second choice with "28,26,12", the longest side is 28, so... a = 12 b = 26 c = 28 a^2 + b^2 = c^2 12^2 + 26^2 = 28^2 144 + 676 = 784 820 = 784 Since that last equation is false, this means that these side lengths do NOT form a right triangle So choice B is out

OpenStudy (waleed_imtiaz):

Option A is a confirmed right triangle..... Use pythagoras theorem.... c^2=a^2+b^2 here.... a=198.b=40....and c should be 202 =39204+1600 c^2= 40804 take square root... c=202